# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v1

@@ 5.5 Parameter Sensitivity
To address RQ4, we carry out parameter sensitivity experiments on the WikiPeople-4 and JF17K-3 datasets, including: a) effect of label smoothing; b) effect of embedding dimension $d$; c) effect of temperature $\gamma$, and d) effect of initialization range. The corresponding results are shown in Figure 4.
@@
Effect of Label Smoothing. Label smoothing is widely used in tasks related to knowledge graphs [52, 53, 54]. It introduces a degree of uncertainty, making the model more adaptable to different data distributions during the training process and further suppresses model overconfidence and improves its generalization ability and robustness. From Figure 4(a), we observe that the values of label smoothing have an impact on the performance, but there is no evident pattern. We need to make appropriate adjustments to its values based on the dataset. For instance, we select 0.8 and 0.5 as the most suitable label smoothing values for the WikiPeople-4 and JF17K-3 datasets, respectively. A smoothing value that is too small or too large may cause performance degradation. A smoothing value that is too large may cause the model to overly believe in the smoothed label distribution, thereby reducing the model’s fitting performance on the training set. A smoothing value that is too small may not effectively prevent the model from overfitting the training data, causing the model to be overconfident in the labels of the training samples.
@@
5.6 Complexity Analysis
To address RQ5, we analyze HyperMono’s complexity on the WD50K, WikiPeople, and JF17K datasets from two perspectives: time complexity and space complexity. The corresponding results are shown in Table VIII. The time complexity is measured by the number of floating-point operations (#FLOPs) required during training phase, while the space complexity refers to the amount of a model’s parameters (#Params). The larger the values of #FLOPs and #Params are, the more computing power and higher memory usage are required during the training process. We observe that compared with HyperFormer, HyperMono requires less time and space overhead in each dataset. Specifically, on the WD50K dataset, HyperMono only requires nearly half the #FLOPs and one-sixth the #Params of HyperFormer, which demonstrates that HyperMono is more efficient and can achieve better results while occupying less time and space.
TABLE VIII: Amount of parameters and calculation of the HyperFormer and HyperMono models.
Metrics Model WD50K WikiPeople JF17K
#FLOPs HyperFormer 118.397G 118.167G 118.070G
HyperMono 69.600G 69.371G 69.273G
#Params HyperFormer 66.956M 66.956M 66.956M
HyperMono 10.645M 10.645M 10.645M
6 Conclusions
In this paper, we propose HyperMono, a framework for hyper-relational graph completion that properly implements two-stage reasoning and faithfully captures qualifier monotonicity. Specifically, we model each qualifier as a shrinking of the main triple cone to a qualifier cone to realize qualifier monotonicity. At the same time, we facilitate the integration of coarse-grained inference results derived solely from main triples and fine-grained inference results obtained from hyper-relational facts with qualifiers. Empirical experiments under various settings on the WD50K, WikiPeople, and JF17K datasets demonstrate HyperMono’s robust performance, when compared to competitive baselines. Additionally, we present various ablation studies on model transferability, parameter sensitivity, and complexity analysis.
7 Future Work
As for future work, on the one hand, we believe that it is necessary to pay special attention to numerical attribute knowledge in hyper-relational knowledge. It is unscientific to simply treat numbers as an independent entity. Therefore, we plant to construct a hyper-relational knowledge graph with discrete numbers for further research. On the other hand, there is some readily available schema knowledge for the entities in the knowledge graph. How to elegantly integrate this knowledge into the representation process is also an interesting line of research.

### v2

@@ D   Parameter Sensitivity
To address RQ3, we carry out parameter sensitivity experiments on the WikiPeople-3, WikiPeople-4, JF17K-3, and JF17K-4 datasets, including: a) effect of label smoothing; b) effect of embedding dimension $d$; c) effect of temperature $\gamma$, and d) effect of initialization range. The corresponding results are shown in Figure 5.
@@
Effect of Label Smoothing. Label smoothing is widely used in tasks related to knowledge graphs (Zhang et al., 2020; Liu et al., 2022b, a). It introduces a degree of uncertainty, making the model more adaptable to different data distributions during the training process. From Figure 5(a), we observe that the values of label smoothing have an impact on the performance, but there is no evident pattern. We need to make appropriate adjustments to its values based on the dataset. For instance, we select 0.8 and 0.5 as the most suitable label smoothing values for the WikiPeople-4 and JF17K-3 datasets, respectively.
Table 7. Evaluation of the knowledge hypergraph completion task in the mixed arity setting on the N-WikiPeople, N-JF17K and N-FB-AUTO datasets. Best scores are highlighted in bold, the second best scores are highlighted in normal.
Methods N-WikiPeople N-JF17K N-FB-AUTO
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
RAE (Zhang et al., 2018) 0.253 0.118 0.343 0.463 0.396 0.312 0.433 0.561 0.703 0.614 0.764 0.854
NaLP (Guan et al., 2019) 0.338 0.272 0.364 0.466 0.310 0.239 0.334 0.450 0.672 0.611 0.712 0.774
HINGE (Rosso et al., 2020) 0.333 0.259 0.361 0.477 0.473 0.397 0.490 0.618 0.678 0.630 0.706 0.765
NeuInfer (Guan et al., 2020) 0.350 0.282 0.381 0.467 0.451 0.373 0.484 0.604 0.737 0.700 0.755 0.805
HypE (Fatemi et al., 2020) 0.263 0.127 0.355 0.486 0.494 0.399 0.532 0.650 0.804 0.774 0.824 0.856
RAM (Liu et al., 2021) 0.380 0.279 0.455 0.539 0.539 0.463 0.573 0.690 0.830 0.803 0.851 0.876
HyperMLN (Chen et al., 2022) 0.351 0.270 0.394 0.497 0.556 0.482 0.597 0.717 0.831 0.803 0.851 0.877
tNaLP+ (Guan et al., 2023) 0.339 0.269 0.369 0.473 0.449 0.370 0.484 0.598 0.729 0.645 0.748 0.826
S2S (Di et al., 2021) 0.372 0.277 0.439 0.533 0.528 0.457 0.570 0.690 - - - -
HyConvE (Wang et al., 2023) 0.362 0.275 0.388 0.501 0.590 0.478 0.610 0.729 0.847 0.820 0.872 0.901
MSeaKG (Di and Chen, 2023) 0.392 0.290 0.468 0.553 0.561 0.475 0.591 0.705 - - - -
HJE (Li et al., 2024b) 0.450 0.375 0.487 0.582 0.590 0.507 0.613 0.729 0.872 0.848 0.886 0.903
PolygonE (Yan et al., 2022) 0.431 0.334 0.454 0.568 0.565 0.485 0.602 0.708 0.858 0.826 0.871 0.921
HyCubE (Li et al., 2024a) 0.438 0.358 0.482 0.575 0.581 0.507 0.615 0.731 0.884 0.863 0.897 0.920
@@
E   Model Transferability
To address RQ5, we further validate the transfer ability of HyperMono. To this aim, we also conduct experiments on the Knowledge Hypergraph Completion (KHC) task. We start by introducing the task definition, dataset descriptions, and baselines. We only introduce necessary background, for a detailed discussion, interested readers are directed to the relevant literature (Di et al., 2021; Wang et al., 2023; Di and Chen, 2023; Li et al., 2024b; Yan et al., 2022; Li et al., 2024a).
Datasets. We test the performance of HyperMono on the KHC task. We conduct experiments on hypergraphs with mixed arity: N-WikiPeople, N-JF17K, and N-FB-AUTO and with fixed arity: N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3, and N-JF17K-4. The statistics of the corresponding datasets are consistent with existing literatures (Di et al., 2021; Wang et al., 2023; Di and Chen, 2023; Li et al., 2024b; Yan et al., 2022; Li et al., 2024a).
Baselines. For knowledge hypergraph completion task, we select 14 state-of-the-art baselines, including RAE (Zhang et al., 2018), NaLP (Guan et al., 2019), HINGE (Rosso et al., 2020), NeuInfer (Guan et al., 2020), HypE (Fatemi et al., 2020), RAM (Liu et al., 2021), HyperMLN (Chen et al., 2022), tNaLP+ (Guan et al., 2023), S2S (Di et al., 2021), HyConvE (Wang et al., 2023), MSeaKG (Di and Chen, 2023), HJE (Li et al., 2024b), PolygonE (Yan et al., 2022), and HyCubE (Li et al., 2024a).
Table 8. Evaluation of the knowledge hypergraph completion task in the fixed arity setting on the N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3 and N-JF17K-4 datasets. Best scores are highlighted in bold, the second best scores are highlighted in normal.
Methods N-WikiPeople-3 N-WikiPeople-4 N-JF17K-3 N-JF17K-4
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
RAE (Zhang et al., 2018) 0.239 0.168 0.252 0.379 0.150 0.080 0.149 0.273 0.505 0.430 0.532 0.644 0.707 0.636 0.751 0.835
NaLP (Guan et al., 2019) 0.301 0.226 0.327 0.445 0.342 0.237 0.400 0.540 0.515 0.431 0.552 0.679 0.719 0.673 0.742 0.805
HINGE (Rosso et al., 2020) 0.338 0.255 0.360 0.508 0.352 0.241 0.419 0.557 0.587 0.509 0.621 0.738 0.745 0.700 0.775 0.842
NeuInfer (Guan et al., 2020) 0.355 0.262 0.388 0.521 0.361 0.255 0.424 0.566 0.622 0.533 0.658 0.770 0.765 0.722 0.808 0.871
HypE (Fatemi et al., 2020) 0.266 0.183 - 0.443 0.304 0.191 - 0.527 0.364 0.255 - 0.573 0.408 0.300 - 0.627
RAM (Liu et al., 2021) 0.254 0.190 - 0.383 0.226 0.161 - 0.367 0.578 0.505 - 0.722 0.743 0.701 - 0.845
HyperMLN (Chen et al., 2022) 0.252 0.193 - 0.385 0.224 0.167 - 0.370 0.574 0.501 - 0.711 0.734 0.687 - 0.831
tNaLP+ (Guan et al., 2023) 0.270 0.185 - 0.444 0.344 0.223 - 0.578 0.411 0.325 - 0.617 0.630 0.531 - 0.722
S2S (Di et al., 2021) 0.386 0.299 0.421 0.559 0.391 0.270 0.470 0.600 0.740 0.676 0.770 0.860 0.822 0.761 0.853 0.924
HyConvE (Wang et al., 2023) 0.318 0.240 - 0.482 0.386 0.271 - 0.607 0.729 0.670 - 0.861 0.827 0.770 - 0.931
MSeaKG (Di and Chen, 2023) 0.403 - - 0.579 0.409 - - 0.624 0.754 - - 0.889 0.833 - - 0.938
HJE (Li et al., 2024b) 0.381 0.303 - 0.596 0.388 0.276 - 0.607 0.707 0.637 - 0.851 0.827 0.775 - 0.928
HyCubE (Li et al., 2024a) 0.322 0.247 - 0.469 0.370 0.252 - 0.583 0.615 0.544 - 0.752 0.792 0.744 - 0.879
Performance Analysis. The results of the knowledge hypergraph completion experiments in the mixed arity and fixed arity settings are respectively shown in Table 7 and Table 8. We observe that our HyperMono consistently achieves state-of-the-art performance on all benchmark datasets. For the mixed arity scenario, compared with the best-performing baseline HyCubE, HyperMono obtains improvements of 4.3% on N-WikiPeople and 9.5% on N-JF17K in the MRR metric. On the N-FB-AUTO dataset, HyperMono virtually matches the results of the best performing baseline. These results demonstrate the capability and effectiveness of HyperMono for capturing and extracting various types of information and knowledge. The same phenomenon occurs in the fixed arity scenario, HyperMono respectively obtains performance improvements of 13.2%, 11.6%, 17.8%, and 9.6% in the MRR metric on the N-WikiPeople-3, N-WikiPeople-4, N-JF17K-3, and N-JF17K-4 datasets, when compared with the best baseline HyCubE. Such a notable performance improvement is mainly due to our explicit treatment of qualifier monotonicity during the encoding process. This feature is more suitable for scenarios where arity is fixed, because introducing qualifier monotonicity can greatly reduce the scope of the optional answer space to obtain the final answer.